OpenStudy (anonymous):

Trains A and B are traveling in the same direction on parallel tracks. Train A is traveling at 60 mph and train B is traveling 80 mph. Train A passes a station at 2:15 am. If train B passes the same station at 2:30 am What time will train B catch up to train A.

OpenStudy (amistre64):

they are 15 minutes apart; what is the head start for train a then? 60(1/4)

OpenStudy (amistre64):

when 60t + headstart = 80t is the time they are equal

OpenStudy (amistre64):

headstart = t(80-60) t = headstart/20 then

OpenStudy (amistre64):

headstart = 15 miles 15/20 = hours

OpenStudy (anonymous):

SO how do I go about figuring out the exact time?

OpenStudy (amistre64):

well, since train B takes 45 minutes to catch up to train A from the time it leaves the stateion at 2:30; id take a gander and say: 2:30 + :45

OpenStudy (anonymous):

ok that makes sense!

OpenStudy (anonymous):

Thank you!